Suppose that 181.1 J of heat energy is required to completely vaporize 3.9 g of a pure liquid. If the molar mass of the liquid is 128.34 g/mol, what is the molar enthalpy of vaporization of the liquid?


1
Expert's answer
2022-01-28T07:36:02-0500


Moles evaporated"=\frac{3.9}{128.34}"


Molar enthalpy of evaporation"=\frac{128.34}{3.9}×\frac{181.1}{1000}"


"=5.9596kJ\/mol"